# What Expressivity Theory Misses: Message Passing Complexity for GNNs

### v1

@@ Appendix A Extended Related Work
While iso expressivity theory is task-agnostic, recent works have proposed studying specific graph-related tasks to get a more fine-grained hierarchy of the expressivity of higher-order MPNNs (beyond the WL test). For example, expressivity has been analyzed in the context of substructure recognition Zhang et al. [2024] and graph biconnectivity Zhang et al. [2023].
A separate line of work characterizes the node-level functions computable by MPNNs in terms of fragments of first-order logic Grohe [2024]: Barceló et al. [2020] showed that standard MPNNs with sum aggregation and ReLU activations can uniformly express graded modal logic. Subsequent work analyzed the effect of different activation functions Khalife and Tonelli-Cueto [2024], and more recently, these logical characterizations were extended to various higher-order MPNNs Xu [2025].
However, all these works only provide existence results: they show whether there exists some model instance of an architecture that can solve a task or not. Such results neglect practical limitations like information loss (see in phenomenons like over-squashing or over-smoothing) occurring in trained real-world models. Moreover, the characterizations are inherently binary (solvable or not), offering limited insight into the practical difficulty.
@@
Some works explicitly account for the lossy information propagation inherent in practical GNNs. Negative results have been derived by considering limited hidden dimensions Loukas [2020] or by analyzing the maximal "mixing" of node representations Giovanni et al. [2024]. However, these approaches are restricted to specific model architectures and provide binary results, rather than a nuanced understanding of how lossy propagation impacts practical performance.
@@
To address the binary nature of classical expressivity theory, some works define distances between graphs to represent the difficulty of distinguishing them. Examples include tree-based distances Chuang and Jegelka [2022], graphon-based distances Böker et al. [2023], and Wasserstein-based distances Chen et al. [2022]. While these metrics provide a continuous measure, they remain task-agnostic and are independent of model architecture.
@@
Classical expressivity theory evaluates GNNs on the set of all possible graphs, which may not align with the restricted graph families encountered in real-world applications. Recent works examine expressivity within specific graph families, such as planar graphs Dimitrov et al. [2023] and outer-planar graphs Bause et al. [2025], providing more relevant insights for certain domains.
@@
A plethora of higher-order MPNNs Bodnar et al. [2021a, b], Wollschläger et al. [2024], Morris et al. [2019] have been developed to surpass the WL test in iso expressivity. To quantify and compare their expressivity corresponding (higher-order) variants of the WL test Morris et al. [2019], Bodnar et al. [2021a], Wollschläger et al. [2024], Wang et al. [2022] have been developed. Besides iso expressivity, they are analyzed by their ability to recognize substructures Zhang et al. [2024] or solve the graph biconnectivity problem Zhang et al. [2023]. However, like iso expressivity, these methods remain binary and do not account for practical limitations of MPNNs like over-squashing. Furthermore, we show in Tab.˜1 similar to Zopf [2022] that the standard WL test already suffices to distinguish almost all graphs in standard benchmarks. Hence, expressivity theory statements that only focus on expressivity beyond the WL test cannot offer explanations for performance differences of models that are at least WL expressive on these popular benchmarks.
@@
Several fundamental limitations of GNNs have been identified in practical applications. Under-reaching Barceló et al. [2020] occurs when nodes cannot access information from distant parts of the graph due to insufficient layer depth, effectively limiting the receptive field of each node. Over-squashing Alon and Yahav [2021], Topping et al. [2022], Giovanni et al. [2023] describes how graph topology can create bottlenecks in message passing, preventing effective information flow between nodes even when they are theoretically within each other’s receptive field. This is distinct from over-smoothing Li et al. [2018], where increasing the number of layers causes node representations to become indistinguishable as all nodes converge to the same representation. These limitations highlight a crucial gap between theoretical expressivity and practical capabilities of GNNs.
@@
A few other works highlight a disconnect between theoretical expressivity limitations and the practical capabilities of standard GNNs. While expressivity theory establishes that standard GNNs cannot count any non-trivial substructure Chen et al. [2020], recent studies reveal important exceptions. Pellizzoni et al. [2025] identify specific conditions on the graph family $\mathcal{G}$} under which standard GNNs can count substructures and demonstrate that many real-world datasets satisfy these conditions for important substructures like small cycles. Similarly, Kanatsoulis and Ribeiro [2024] showed that adding randomized node features enables standard GNNs to count small cycles.
@@ Appendix B Extending MPC: More Architectures, Graph-Level Functions, and Feature Noise
In the following, we show how MPC can be extended to a wide range of architectures beyond standard MPNNs and to graph-level functions. Additionally, we sketch a potential extension to include additional sources of complexity, such as feature noise.
@@ B.1 General Message Passing Framework
The transformation $t$ defines the MP graph structure $\tilde{{G}}$} by determining which nodes exchange messages. While standard MPNNs used the input graph G{G} directly as $\tilde{{G}}$}, recent architectures introduce modifications like virtual nodes Gilmer et al. [2017], rewired edges Topping et al. [2022], or higher-order graphs Wollschläger et al. [2024]. We assume $t$ preserves original nodes $\mathcal{V}_{G}$} while potentially adding nodes or adding/modifying edges. A general MPNN architecture $\mathcal{M}$} is defined through the transformation $t$ and the number of layers $L$, characterizing which nodes exchange information. It abstracts away from the specific choice of aggregation method and update function, which is a sensible simplification since many recent models treat the choice of aggregation method as a hyperparameter and use MLPs as update functions (providing universal function approximation capabilities).
@@
Standard MPNNs Standard architectures like GCN Kipf and Welling [2017], GraphSage Hamilton et al. [2017], GIN Xu et al. [2019] perform message passing directly on the input graph, i.e., ${\tilde{{G}}}=t({G})={G}$. They differ in aggregation method (mean, sum, degree-normalized mean). Some architectures, like GCN, restrict the possible choices of msg and upd function and do not distinguish self-loops from normal edges, i.e., ${\texttt{msg}}_{0}={\texttt{msg}}_{1}$. Note that, as discussed above, our MPNN framework abstracts away these architectural choices.
Virtual nodes, rewiring and additional features Modifications to the input graph such as additional virtual nodes Gilmer et al. [2017], edge rewiring Topping et al. [2022], or additional node features (e.g., including substructure information like in GSN Bouritsas et al. [2023]) can be directly modeled in our framework by choosing the appropriate transformation function $t$.
Higher-order MPNNs Higher-order MPNNs that augment the input graph with higher-level structures are also representable. FragNet Wollschläger et al. [2024] builds an additional higher-level graph of fragments (for an arbitrary fragmentation scheme). We will focus here on the FragNet model with a fragmentation scheme identifying rings without edge representations. We model this by adding additional fragment nodes connected to all constituent input nodes, with labels corresponding to fragment types. The CIN Bodnar et al. [2021a] model follows a similar approach by having representations for all original nodes, edges, and CW-cells (rings). Upper messages are messages from nodes to their corresponding edge representation, and from edges to the corresponding ring representation (if any). Boundary messages are messages between nodes sharing an edge and between edges that are both in the same ring. The initial representation of CW cells is an aggregation of the features of their constituent nodes. We represent edges and CW cells (rings) by additional nodes in $\tilde{{G}}$}. The initial feature of a ring node is the aggregation of all node features of the ring. We have edges between: neighboring original nodes, original nodes and their edge nodes, edge nodes within the same ring, and edge nodes and their rings (if any). Note that we cannot represent in our framework that boundary messages also contain information from the corresponding upper neighborhood. Importantly, both fragNet and CIN use different update functions for different types of messages. This too can not be modeled in our formulation of $\mathcal{M}$}; instead, we assume just one update function. While one could, in theory, extend our framework to include different kinds of message updates, we choose against this as our existing framework, with this simplification, already captures the models’ empirical performance well.
@@
$\textup{WLC}_{\mathcal{M}}(f_{v},{G})=\begin{cases}\infty&\text{if }{\text{WL}}^{L}_{v}(t({G}))\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})\\ 0&\text{else}\end{cases}$
@@ Appendix C Proofs and Extended Theorems
Extended Notation Let $\circ$ denote function composition. Additionally, let $\alpha^{-1}$ be the preimage of a function $\alpha:X\to Y$, i.e., $\alpha^{-1}(y)=\{x\in X\mid\alpha(x)=y\}$. If $X$ is clear from the context, we will simply write $\alpha\vDash\beta$ instead of $\alpha\vDash_{X}\beta$. We define ${\text{WL}}_{v}:={\text{WL}}^{L}_{v}$ and $\textup{lossyWL}_{v}:=\textup{lossyWL}_{v}^{L}$, i.e., we omit the layer superscript for the final layer $L$. Additionally, we say a message $m_{u\to v}^{l}$ in the lossyWL test is lost, if $Z_{uv}^{l}=0$, otherwise we say it was successful. Let ${M}_{v}({G})$ denote the final output $h_{v}^{L}$ of a model at node $v$. Additionally, we will say a set $\mathcal{S}\subseteq 2^{U}$ of sets over an universe $U$ is upward-closed if $A\in\mathcal{S}$ implies that any $B\subseteq U$ with $A\subseteq B$ is also in $\mathcal{S}$, i.e., $B\in\mathcal{S}$. $\mathbf{1}[\text{condition}]$ denotes the indicator function defined as $1$ if the condition is true and 0 else. We use ring and cycle interchangeably.
@@ C.2 Connection of standard MPNNs to WL
There exists a model $M\in{{\mathcal{M}}_{S}}$ that performs exactly the WL test, i.e., for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ $M_{v}({G})={\text{WL}}_{v}({G})$.
@@
We show this inductively over the number of layers $L$. For $L=0$ this follows immediately from ${\text{WL}}^{0}_{v}({G})={\bm{X}}_{v}$ and $M^{0}_{v}({G})={\bm{X}}_{v}$ for any 0-layer model ${M}^{0}\in{\mathcal{M}}$. For $L=l>0$, assume that there exists a model $M^{l-1}$ with $M^{l-1}_{v}={\text{WL}}_{v}^{l-1}$. Then consider a model $M^{l}$ that executes $M^{l-1}$ in the first l1l-1 layers:
$m_{w\to v}^{l}:=\begin{cases}{\texttt{msg}}_{0}^{l}({\text{WL}}_{w}^{l-1})&\text{if}\>\>w=v\\ {\texttt{msg}}_{1}^{l}({\text{WL}}_{w}^{l-1})&\text{else}\end{cases}\quad\text{and}\quad h_{v}^{l}:={\texttt{upd}}^{l}({\texttt{agg}}\left(\{\mskip-5.0mu\{m_{w\to v}^{l}\mid w\in{N}_{G}(v)\cup\{v\}\}\mskip-5.0mu\}\right))$
@@
$h_{v}^{l}={\text{WL}}_{v}^{l}=\operatorname{HASH}\left({\text{WL}}^{l-1}_{v},\{\mskip-5.0mu\{{\text{WL}}^{l-1}_{u}\mid u\in{N}_{G}(v)\}\mskip-5.0mu\}\right).$
@@
${\text{WL}}_{v}^{l}({G})$ $=\operatorname{HASH}\left({\text{WL}}^{l-1}_{v},\{\mskip-5.0mu\{{\text{WL}}^{l-1}_{u}\mid u\in{N}_{G}(v)\}\mskip-5.0mu\}\right)$
$={\texttt{upd}}^{l}\left(\sum_{w\in{N}_{G}(v)\cup\{v\}}{\texttt{msg}}^{l}_{\mathbf{1}[w=v]}({\text{WL}}_{w}^{l-1})\right)$
@@ C.3 When MPC becomes infinite: Connections to iso expressivity theory
${\text{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
We will first assume that ${\text{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})$. The proof idea for this direction is that with positive probability lossyWL loses no to $v$ visible message and performs exactly the normal WL test. This coloring will always be different from lossyWL colors with visible lost messages and can, hence, by assumption, be used to deduce $f_{v}$.
@@
$\textup{lossyWL}_{v}({G},s_{1})=\textup{lossyWL}_{w}({G}_{2},s_{2})\implies{\text{WL}}_{v}({G})={\text{WL}}_{w}({G}_{2})$
@@
$\textup{lossyWL}_{v}({G},s_{1})\vDash_{{{\mathcal{G}}_{\mathcal{V}}}\times S}{\text{WL}}_{v}({G}).$
@@
$\textup{lossyWL}_{v}({G},s_{1})\vDash{\text{WL}}_{v}({G})\vDash f_{v}({G}),$
@@
Now we show that from ${\text{WL}}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})$, it follows that $\mathbb{P}[\textup{lossyWL}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})]=0$, i.e., losing messages cannot make the probabilistic WL test more expressive than the deterministic WL test. ${\text{WL}}_{v}({G})\nvDash f_{v}({G})$ implies there is ${G}^{\prime},w\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\text{WL}}_{v}({G})={\text{WL}}_{w}({G}^{\prime})$ but $f_{v}({G})\neq f_{w}({G}^{\prime})$. Then, for every seed $s_{1}\in S$, there exists a seed $s_{2}\in S$ such that
@@
${\text{WL}}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
${\text{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
If there exists ${M}\in{\mathcal{M}}$ such that ${M}_{v}({G})=f_{v}({G})$ for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$, it follows directly from Footnote˜1 that there does not exist ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\textup{MPC}}_{{\mathcal{M}}_{S}}(f_{v},{G})=\infty$.
@@
${\text{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
${\text{WL}}\vDash_{{\mathcal{G}}_{\mathcal{V}}}f.$
@@
If ${\textup{MPC}}{}_{\mathcal{M}}(\alpha,{G})=\infty$ for some ${G}\in{\mathcal{G}}^{*}$, then by Footnote˜1, there exists ${G}^{\prime}\in{\mathcal{G}}^{*}$ such that ${M}({G})={M}({G}^{\prime})$ for all ${M}\in{\mathcal{M}}$ but $\alpha({G})\neq\alpha({G}^{\prime})$. From this follows directly that $\mathcal{M}$} is not at least as expressive as $\alpha$.
Now assume that there does not exist ${G}\in{\mathcal{G}}^{*}$ with ${\textup{MPC}}{}_{\mathcal{M}}(\alpha,{G})=\infty$. Then by Footnote˜1 there exists $M\in{\mathcal{M}}$ with $M({G})=\alpha({G})$ for all ${G}\in{\mathcal{G}}^{*}$. Therefore, $\mathcal{M}$} is at least as expressive as $\alpha$. ∎

### v2

@@ Appendix A Extended Related Work
While iso expressivity theory is task-agnostic, recent works have proposed studying specific graph-related tasks to get a more fine-grained hierarchy of the expressivity of higher-order MPNNs (beyond the WL test). For example, expressivity has been analyzed in the context of substructure recognition [53] and graph biconnectivity [52].
A separate line of work characterizes the node-level functions computable by MPNNs in terms of fragments of first-order logic [23]: Barceló et al. [3] showed that standard MPNNs with sum aggregation and ReLU activations can uniformly express graded modal logic. Subsequent work analyzed the effect of different activation functions [29], and more recently, these logical characterizations were extended to various higher-order MPNNs [50].
However, all these works only provide existence results: they show whether there exists some model instance of an architecture that can solve a task or not. Such results neglect practical limitations like information loss (seem in phenomenons like over-squashing or over-smoothing) occurring in trained real-world models. Moreover, the characterizations are inherently binary (solvable or not), offering limited insight into the practical difficulty.
@@
Some works explicitly account for the lossy information propagation inherent in practical GNNs. Negative results have been derived by considering limited hidden dimensions [35] or by analyzing the maximal "mixing" of node representations [20]. However, these approaches are restricted to specific model architectures and provide binary results, rather than a nuanced understanding of how lossy propagation impacts practical performance in arbitrary tasks.
@@
To address the binary nature of classical expressivity theory, some works define distances between graphs to represent the difficulty of distinguishing them. Examples include tree-based distances [12], graphon-based distances [7], and Wasserstein-based distances [10]. While these metrics provide a continuous measure, they remain task-agnostic and are independent of model architecture.
@@
Classical expressivity theory evaluates GNNs on the set of all possible graphs, which may not align with the restricted graph families encountered in real-world applications. Recent works examine expressivity within specific graph families, such as planar graphs [13] and outer-planar graphs [4], providing more relevant insights for certain domains.
@@
A plethora of higher-order MPNNs [5, 6, 46, 36] have been developed to surpass the WL test in iso expressivity. To quantify and compare their expressivity corresponding (higher-order) variants of the WL test [36, 5, 46, 43] have been developed. However, we show in Tab.˜1 similar to Zopf [54] that the standard WL test already suffices to distinguish almost all graphs in standard benchmarks. Hence, expressivity theory statements that only focus on expressivity beyond the WL test cannot offer explanations for performance differences of models that are at least WL expressive on these popular benchmarks.
@@
Several fundamental limitations of GNNs have been identified in practical applications. Under-reaching [3] occurs when nodes cannot access information from distant parts of the graph due to insufficient layer depth, effectively limiting the receptive field of each node. Over-squashing [1, 41, 19] describes how graph topology can create bottlenecks in message passing, preventing effective information flow between nodes even when they are theoretically within each other’s receptive field. This is distinct from over-smoothing [33], where increasing the number of layers causes node representations to become indistinguishable as all nodes converge to the same representation. These limitations highlight a crucial gap between theoretical expressivity and practical capabilities of GNNs.
@@
A few other works highlight a disconnect between theoretical expressivity limitations and the practical capabilities of standard GNNs. While expressivity theory establishes that standard GNNs cannot count any non-trivial substructure [11], recent studies reveal important exceptions. Pellizzoni et al. [40] identify specific conditions on the graph family $\mathcal{G}$} under which standard GNNs can count substructures and demonstrate that many real-world datasets satisfy these conditions for important substructures like small cycles. Similarly, Kanatsoulis and Ribeiro [28] showed that adding randomized node features enables standard GNNs to count small cycles.
@@ Appendix B Extending MPC: More Architectures, Graph-Level Functions, and Feature Noise
In the following, we show how MPC can be extended to a wide range of architectures beyond standard MPNNs and to graph-level functions. Additionally, we sketch a potential extension to include additional sources of complexity, such as feature noise. Furthermore, we sketch how one could easily adapt the framework to operate on temporal graphs.
@@ B.1 General Message Passing Framework
The transformation $t$ defines the MP graph structure $\tilde{{G}}$} by determining which nodes exchange messages. While standard MPNNs used the input graph G{G} directly as $\tilde{{G}}$}, recent architectures introduce modifications like virtual nodes [18], rewired edges [41], or higher-order graphs [46]. We assume $t$ preserves original nodes $\mathcal{V}_{G}$} while potentially adding nodes or adding/modifying edges. A general MPNN architecture $\mathcal{M}$} is defined through the transformation $t$ and the number of layers $L$, characterizing which nodes exchange information. It abstracts away from the specific choice of aggregation method and update function, which is a sensible simplification since many recent models treat the choice of aggregation method as a hyperparameter and use MLPs as update functions (providing universal function approximation capabilities).
@@
Standard MPNNs Standard architectures like GCN [31], GraphSage [25], GIN [48] perform message passing directly on the input graph, i.e., ${\tilde{{G}}}=t({G})={G}$. They differ in aggregation method (mean, sum, degree-normalized mean). Some architectures, like GCN, restrict the possible choices of msg and upd function and do not distinguish self-loops from normal edges, i.e., ${\texttt{msg}}_{0}={\texttt{msg}}_{1}$. Note that, as discussed above, our MPNN framework abstracts away these architectural choices.
Virtual nodes, rewiring and additional features Modifications to the input graph such as additional virtual nodes [18], edge rewiring [41], or additional node features (e.g., including substructure information like in GSN [8]) can be directly modeled in our framework by choosing the appropriate transformation function $t$.
Higher-order MPNNs Higher-order MPNNs that augment the input graph with higher-level structures are also representable. FragNet [46] builds an additional higher-level graph of fragments (for an arbitrary fragmentation scheme). We will focus here on the FragNet model with a fragmentation scheme identifying rings without edge representations. We model this by adding additional fragment nodes connected to all constituent input nodes, with labels corresponding to fragment types. The CIN [5] model follows a similar approach by having representations for all original nodes, edges, and CW-cells (rings). Upper messages are messages from nodes to their corresponding edge representation, and from edges to the corresponding ring representation (if any). Boundary messages are messages between nodes sharing an edge and between edges that are both in the same ring. The initial representation of CW cells is an aggregation of the features of their constituent nodes. We represent edges and CW cells (rings) by additional nodes in $\tilde{{G}}$}. The initial feature of a ring node is the aggregation of all node features of the ring. We have edges between: neighboring original nodes, original nodes and their edge nodes, edge nodes within the same ring, and edge nodes and their rings (if any). Note that we cannot represent in our framework that boundary messages also contain information from the corresponding upper neighborhood. Importantly, both fragNet and CIN use different update functions for different types of messages. This too can not be modeled in our formulation of $\mathcal{M}$}; instead, we assume just one update function. While one could, in theory, extend our framework to include different kinds of message updates, we choose against this as our existing framework, with this simplification, already captures the models’ empirical performance well.
@@
$\textup{WLC}_{\mathcal{M}}(f_{v},{G})=\begin{cases}\infty&\text{if }{\textup{WL}}^{L}_{v}(t({G}))\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})\\ 0&\text{else}\end{cases}$
@@ B.3 Additional Sources of Complexity: Feature Noise
B.4 Temporal Graphs
MPC can be naturally extended to temporal graphs as well. Consider snapshot graphs ${G}^{\tau}$ at time point $\tau$, as defined in Longa et al. [34]. The complexity for a task $f_{v}$ then becomes time dependent:
${\textup{MPC}}_{\mathcal{M}}^{\tau}(f_{v},G):=-\log\mathbb{P}[\textup{lossyWL}_{v}^{L}(t({G}^{\tau}))\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}(G)].$
This extension opens up interesting directions for future work, such as analyzing how the complexity of specific tasks evolves over time as the underlying graph changes.
@@ Appendix C Proofs and Extended Theorems
Extended Notation Let $\circ$ denote function composition. Additionally, let $\alpha^{-1}$ be the preimage of a function $\alpha:X\to Y$, i.e., $\alpha^{-1}(y)=\{x\in X\mid\alpha(x)=y\}$. If $X$ is clear from the context, we will simply write $\alpha\vDash\beta$ instead of $\alpha\vDash_{X}\beta$. We define ${\textup{WL}}_{v}:={\textup{WL}}^{L}_{v}$ and $\textup{lossyWL}_{v}:=\textup{lossyWL}_{v}^{L}$, i.e., we omit the layer superscript for the final layer $L$. Additionally, we say a message $m_{u\to v}^{l}$ in the lossyWL test is lost, if $Z_{uv}^{l}=0$, otherwise we say it was successful. Let ${M}_{v}({G})$ denote the final output $h_{v}^{L}$ of a model at node $v$. Additionally, we will say a set $\mathcal{S}\subseteq 2^{U}$ of sets over an universe $U$ is upward-closed if $A\in\mathcal{S}$ implies that any $B\subseteq U$ with $A\subseteq B$ is also in $\mathcal{S}$, i.e., $B\in\mathcal{S}$. $\mathbf{1}[\text{condition}]$ denotes the indicator function defined as $1$ if the condition is true and 0 else. We use the terms "ring" and "cycle" interchangeably.
@@ C.2 Connection of standard MPNNs to WL
There exists a model $M\in{{\mathcal{M}}_{S}}$ that performs exactly the WL test, i.e., for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ $M_{v}({G})={\textup{WL}}_{v}({G})$.
@@
We show this inductively over the number of layers $L$. For $L=0$ this follows immediately from ${\textup{WL}}^{0}_{v}({G})={\bm{X}}_{v}$ and $M^{0}_{v}({G})={\bm{X}}_{v}$ for any 0-layer model ${M}^{0}\in{\mathcal{M}}$. For $L=l>0$, assume that there exists a model $M^{l-1}$ with $M^{l-1}_{v}={\textup{WL}}_{v}^{l-1}$. Then consider a model $M^{l}$ that executes $M^{l-1}$ in the first l1l-1 layers:
$m_{w\to v}^{l}:=\begin{cases}{\texttt{msg}}_{0}^{l}({\textup{WL}}_{w}^{l-1})&\text{if}\>\>w=v\\ {\texttt{msg}}_{1}^{l}({\textup{WL}}_{w}^{l-1})&\text{else}\end{cases}\quad\text{and}\quad h_{v}^{l}:={\texttt{upd}}^{l}({\texttt{agg}}\left(\{\mskip-5.0mu\{m_{w\to v}^{l}\mid w\in{N}_{G}(v)\cup\{v\}\}\mskip-5.0mu\}\right))$
@@
$h_{v}^{l}={\textup{WL}}_{v}^{l}=\operatorname{HASH}\left({\textup{WL}}^{l-1}_{v},\{\mskip-5.0mu\{{\textup{WL}}^{l-1}_{u}\mid u\in{N}_{G}(v)\}\mskip-5.0mu\}\right).$
@@
${\textup{WL}}_{v}^{l}({G})$ $=\operatorname{HASH}\left({\textup{WL}}^{l-1}_{v},\{\mskip-5.0mu\{{\textup{WL}}^{l-1}_{u}\mid u\in{N}_{G}(v)\}\mskip-5.0mu\}\right)$
$={\texttt{upd}}^{l}\left(\sum_{w\in{N}_{G}(v)\cup\{v\}}{\texttt{msg}}^{l}_{\mathbf{1}[w=v]}({\textup{WL}}_{w}^{l-1})\right)$
@@ C.3 When MPC becomes infinite: Connections to iso expressivity theory
${\textup{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
We will first assume that ${\textup{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})$. The proof idea for this direction is that with positive probability lossyWL loses no to $v$ visible message and performs exactly the normal WL test. This coloring will always be different from lossyWL colors with visible lost messages and can, hence, by assumption, be used to deduce $f_{v}$.
@@
$\textup{lossyWL}_{v}({G},s_{1})=\textup{lossyWL}_{w}({G}_{2},s_{2})\implies{\textup{WL}}_{v}({G})={\textup{WL}}_{w}({G}_{2})$
@@
$\textup{lossyWL}_{v}({G},s_{1})\vDash_{{{\mathcal{G}}_{\mathcal{V}}}\times S}{\textup{WL}}_{v}({G}).$
@@
$\textup{lossyWL}_{v}({G},s_{1})\vDash{\textup{WL}}_{v}({G})\vDash f_{v}({G}),$
@@
Now we show that from ${\textup{WL}}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})$, it follows that $\mathbb{P}[\textup{lossyWL}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})]=0$, i.e., losing messages cannot make the probabilistic WL test more expressive than the deterministic WL test. ${\textup{WL}}_{v}({G})\nvDash f_{v}({G})$ implies there is ${G}^{\prime},w\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\textup{WL}}_{v}({G})={\textup{WL}}_{w}({G}^{\prime})$ but $f_{v}({G})\neq f_{w}({G}^{\prime})$. Then, for every seed $s_{1}\in S$, there exists a seed $s_{2}\in S$ such that
@@
${\textup{WL}}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
${\textup{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
If there exists ${M}\in{\mathcal{M}}$ such that ${M}_{v}({G})=f_{v}({G})$ for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$, it follows directly from Footnote˜2 that there does not exist ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\textup{MPC}}_{{\mathcal{M}}_{S}}(f_{v},{G})=\infty$.
@@
${\textup{WL}}_{v}({G})\vDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G}).$
@@
${\textup{WL}}\vDash_{{\mathcal{G}}_{\mathcal{V}}}f.$
@@
If ${\textup{MPC}}{}_{\mathcal{M}}(\alpha,{G})=\infty$ for some ${G}\in{\mathcal{G}}^{*}$, then by Footnote˜2, there exists ${G}^{\prime}\in{\mathcal{G}}^{*}$ such that ${M}({G})={M}({G}^{\prime})$ for all ${M}\in{\mathcal{M}}$ but $\alpha({G})\neq\alpha({G}^{\prime})$. From this follows directly that $\mathcal{M}$} is not at least as expressive as $\alpha$.
Now assume that there does not exist ${G}\in{\mathcal{G}}^{*}$ with ${\textup{MPC}}{}_{\mathcal{M}}(\alpha,{G})=\infty$. Then by Footnote˜2 there exists $M\in{\mathcal{M}}$ with $M({G})=\alpha({G})$ for all ${G}\in{\mathcal{G}}^{*}$. Therefore, $\mathcal{M}$} is at least as expressive as $\alpha$. ∎